OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

please help !!

OpenStudy (shamim):

equation of line which go through 2 points is\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{1})\]

OpenStudy (shamim):

here \[x _{1}=0,y _{1}=0,x _{2}=4,y _{2}=-4\]

OpenStudy (shamim):

just put these value in above equation

OpenStudy (shamim):

did u get it

OpenStudy (anonymous):

y- -4 = x-0

OpenStudy (shamim):

no

OpenStudy (shamim):

\[y-0=\frac{ -4-0 }{ 4-0 }(x-0)\]

OpenStudy (anonymous):

oh i pluged them in wrong

OpenStudy (anonymous):

y-0= -4/4(x-0)

OpenStudy (shamim):

ya

OpenStudy (shamim):

then?

OpenStudy (anonymous):

does -4/4 = 0 or -1

OpenStudy (shamim):

-1

OpenStudy (shamim):

anyway\[-4+4=0\]

OpenStudy (shamim):

but\[\frac{ -4 }{ 4 }=-1\]

OpenStudy (shamim):

\[\frac{ 4 }{ 4 }=1\]

OpenStudy (anonymous):

y-0=-1x-(-1)

OpenStudy (shamim):

no

OpenStudy (shamim):

i m showing u every steps

OpenStudy (shamim):

\[y-0=\frac{ -4-0 }{ 4-0 }(x-0)\]

OpenStudy (shamim):

\[y=\frac{ -4 }{ 4 }x\]

OpenStudy (shamim):

\[y=-1x\]

OpenStudy (shamim):

\[y=-x\]

OpenStudy (shamim):

\[y+x=0\]

OpenStudy (shamim):

\[x+y=0\]

OpenStudy (shamim):

feel free to ask any question

OpenStudy (anonymous):

thank you for helping me

OpenStudy (shamim):

welcome

OpenStudy (anonymous):

do you know how to do this Indicate the equation of the given line in standard form. The line containing the diagonal of a square whose vertices are A(-3, 3), B(3, 3), C(3, -3), and D(-3, -3). Find two equations, one for each diagonal.

OpenStudy (shamim):

ya i can solve it

OpenStudy (shamim):

bt u should post ur in ask a question box

OpenStudy (shamim):

anyway i m solving it

OpenStudy (shamim):

so ur 1 line will go through A(-3,3) and C(3,-3)

OpenStudy (anonymous):

im sorry my first time on this

OpenStudy (shamim):

its one diagonal

OpenStudy (shamim):

its ok

OpenStudy (anonymous):

its what

OpenStudy (shamim):

sorry for late

OpenStudy (shamim):

anyway i was solving another math

OpenStudy (shamim):

now ur 1 line will go through A(-3,3) and C(3,-3)

OpenStudy (shamim):

so the formula for ur diagonal line will b\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{1})\]

OpenStudy (shamim):

here \[x _{1}=-3,y _{1}=3,x _{2}=3,y _{2}=-3\]

OpenStudy (shamim):

so\[y-3=\frac{ -3-3 }{ 3-(-3) }(x-(-3))\]

OpenStudy (shamim):

now solve it

OpenStudy (shamim):

did u get it